3. [10 points] Suppose an lDs is 90% accurate, so that there is a 10% chance of false positives or false negatives. There are 500,000 log entries registered by the IDS, where 200 of these entries are actual malicious events. How many times will the alarm be sounded and how many are false negatives?
Answer #1
Predicted Negative Predicted Positive
Negative Cases 499800*90/100=44980 499800*10/100=49980
Positive Cases 200*10/100=20 200*90/100=180

Therefore Aarm will be sounded 49980+180 times (50160 times) and 20 cases are false negative.